OpenStudy (anonymous):

A triangle has sides of legths 4, 3, and 5. Is it a right triangle? Explain.

OpenStudy (lgbasallote):

verify by using pythagorean theorem ;)

OpenStudy (lgbasallote):

\[3^2 + 4^2 = 5^2?\]

Parth (parthkohli):

Try it with the pythagorean theorem: \( \color{Black}{\Rightarrow a^2 + b^2 = c^2 }\) c is the longest side.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3^2 + 4^2 = 5^2 }\) Check the equality

OpenStudy (king):

yeah cuz a^2+b^2=c^2 where a,b<c

OpenStudy (anonymous):

I think it is, they answers given are yes; 4^2 + 3^2 =5 yes, 4^2+3^2 then an equal sign with a line through it. and no with the same date.

OpenStudy (lgbasallote):

same date?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3^2 + 4^2 = 5^2 }\)

Parth (parthkohli):

^ it's not 3^2 + 4^2 = 5

OpenStudy (anonymous):

Same data*

OpenStudy (lgbasallote):

what is 3^2 + 4^2?

Parth (parthkohli):

Oh you mean this sign? - \(\ne\)

OpenStudy (anonymous):

yes that sign

Parth (parthkohli):

First check if \(3^2 + 4^2 = 5^2\)

OpenStudy (lgbasallote):

@RandiHobbs what is 3^2 + 4^2?

OpenStudy (anonymous):

Ummm..7^2?..i dunno =[

OpenStudy (lgbasallote):

no... what is 3^2?

OpenStudy (anonymous):

9?

Parth (parthkohli):

Yep.. now what's 4^2

OpenStudy (lgbasallote):

right! what is 4^2?

OpenStudy (anonymous):

16

OpenStudy (lgbasallote):

what is 9 + 16?

Parth (parthkohli):

yes.. so what's 3^2 + 4^2 supposed to be

OpenStudy (anonymous):

25

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now what is 5^2

Parth (parthkohli):

yes... now 5^2 = ..?

OpenStudy (anonymous):

25

OpenStudy (lgbasallote):

are tehy equal?

Parth (parthkohli):

yes, now is 25 = 25?

Parth (parthkohli):

tehy? the word does not exist

Parth (parthkohli):

:P

OpenStudy (lgbasallote):

lol @ParthKohli how are we saying same things :p

OpenStudy (lgbasallote):

they*

Parth (parthkohli):

lol

OpenStudy (anonymous):

well, i'm not that stupid, so yes, 25=25

OpenStudy (lgbasallote):

then it's a right triangle ^_^

Parth (parthkohli):

lol, if they are equal, then they can be the sides of a right triangle :F

OpenStudy (lgbasallote):

since they are equal

OpenStudy (lgbasallote):

do you get it @RandiHobbs ?

OpenStudy (anonymous):

Then what does the line with the equal sign in it indicate?

OpenStudy (lgbasallote):

\(\ne\) means not equal

OpenStudy (anonymous):

Well, then would a triangle with the side lengths of 6, 8, and 9 be a right triangle?

OpenStudy (lgbasallote):

check.. \[6^2 + 8^2 = 9^2?\]

Parth (parthkohli):

Try it. \(6^2 + 8^2 = 9^2 ?\)

Parth (parthkohli):

lol @lgbasallote

OpenStudy (lgbasallote):

stop copying me >.< :p

OpenStudy (anonymous):

Well 6x6= 36 and 8x8=64 and 9x9= 81

OpenStudy (anonymous):

and 36+64 =100

OpenStudy (lgbasallote):

then 100 \(\ne\) 81

Parth (parthkohli):

Yep

OpenStudy (lgbasallote):

since they're not equal..it's not a right triangle

Parth (parthkohli):

SO they can't be the sides of a right triangle.

OpenStudy (lgbasallote):

ugh seriously?!! :PPP

Parth (parthkohli):

or, you can make the hypotenuse as 10 to get a a right triangle :D

OpenStudy (anonymous):

so it would be obtuse?

Parth (parthkohli):

6,8,10 is a pythagorean triplet cuz 6^2 + 8^2 = 10^2

Parth (parthkohli):

Yeah

Parth (parthkohli):

noo.. it'd be acute

OpenStudy (lgbasallote):

6,8,10 isnt a pythagorean triple...3,4,5 is the pythagorean triple :p lol

OpenStudy (anonymous):

>.<!!!!! you are frying my brain

OpenStudy (lgbasallote):

100 is the side that will make it right 81 < 100 therefore since it is smaller than right...it is acute

OpenStudy (lgbasallote):

100 is the side that will make it a right triangle *

OpenStudy (lgbasallote):

do you get that logic?

Parth (parthkohli):

Obtuse ====> a^2 + b^2 < c^2 Right ====> a^2 + b^2 = c^2 Acute ====> a^2 + b^2 > c^2

Parth (parthkohli):

but in the right triangle, there is a restriction to c.

Parth (parthkohli):

can you tell me what that restriction is :)

OpenStudy (anonymous):

No, ma'am i cannot. =[

Parth (parthkohli):

ma'am? who's ma'am?

OpenStudy (lgbasallote):

well do you at least get the logic i presented @RandiHobbs ?

Parth (parthkohli):

@myininaya lol they think that the drawing's of a girl >.< :P

OpenStudy (anonymous):

I dunno,sir, ma'am. Whichever you prefer. ;D

Parth (parthkohli):

lol

OpenStudy (lgbasallote):

he's good with either ;)

Parth (parthkohli):

'c' in a right triangle is supposed to be the longest side.

Parth (parthkohli):

@lgbasallote lol you.. I'll kill you one day

OpenStudy (lgbasallote):

lol even i was wondering what that rstriction was...it was that all along :p

Parth (parthkohli):

lol the asker isn't even viewing the question >.<

OpenStudy (lgbasallote):

hmm

Parth (parthkohli):

mmh

OpenStudy (lgbasallote):

well even i always go away from my questions...i mean my serious questions

OpenStudy (anonymous):

Alright, new question, my lovely mathmeticians .

OpenStudy (anonymous):

ABC is a right triangle. If AC=4 and BC=10, find AB. leave your answer in simplest radical form. (please explain how to leave your answer in simplest radical form?)

OpenStudy (anonymous):

@lgbasallote @Preetha

OpenStudy (lgbasallote):

could you post this in a new question? this thread is getting laggy